OpenStudy (anonymous):

if "a" is 4 times bigger than "b" and "c" is 2 times smaller than "d", and "b" is "d", find "a". =D

OpenStudy (anonymous):

oh, forgot to say that c is 1

OpenStudy (anonymous):

so I have the following as given: 4A = b c/2 = d b =d and c = 1 right?

OpenStudy (anonymous):

no wait! it's A = 4 b right?

OpenStudy (anonymous):

and c = d/2

OpenStudy (anonymous):

so: c = A/4/2 A = 8 B = 8/4 = 2 D = 2 did I get it right? ^_^

OpenStudy (anonymous):

D = B = 2 :)

OpenStudy (anonymous):

right? .-.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

seriously =-=

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

just created this question freestyle without much thought

OpenStudy (anonymous):

smart one though

OpenStudy (anonymous):

created this question so that you dont have to sit idle

OpenStudy (anonymous):

lol, why can't I sit idle? :)

OpenStudy (anonymous):

because being idle makes deteriorates your brain... oh, btw, do you know a site called SQUIDOO.com ??

OpenStudy (anonymous):

lol, not really, when I'm idling I solve questions from my lectures ^_^ and yes, I've heard of such site, but never visited it though , why?

OpenStudy (anonymous):

because being as smart and knowledgeable as you, you could write articles for squidoo, it is fun

OpenStudy (anonymous):

and you make some money

OpenStudy (anonymous):

so, if you ever think of writing articles online, use squidoo, free, nice, easy, and you can even make some money

OpenStudy (anonymous):

thanks for the advise :)

OpenStudy (anonymous):

oh, did you read my letter?

OpenStudy (anonymous):

You should be a business man since you've got lots of ideas of making money :)

OpenStudy (anonymous):

i agree with you sstarica ... A=8

OpenStudy (anonymous):

^_^